Tommy’s Tile Service is planning on purchasing new tile cleaning equipment that will improve their ability to remove tough stains from ceramic tiles. The company’s contribution margin is 25% and its current break-even point is $487,200 in sales revenue. Purchasing the new equipment will increase fixed costs by $11,000.

Required:

1. Determine the company’s current fixed costs.

2. Determine the company’s new break-even point in sales.

3. After the purchase of the equipment, how much revenue does the company need to generate a profit of $135,000?
Answer #1

Solution:

1. $121,800

2. $531,200

3. $1071,200

Explanation:

1.

Sales needed to break even = Fixed cost / Contribution margin ratio

487200 = Fixed cost / 0.25

Fixed cost = 487200 x 0.25 = $121,800

2.

New fixed costs = 121800 + 11000 = 132,800

Sales needed to break even = 132800 / 0.25 = $531,200

3.

Target sales = (Fixed cost + Target profit) / Contribution margin ratio = (132800 + 135000) / 0.25 =$1071200